# Detecting State Changes in Functional Neuronal Connectivity using Factorial Switching Linear Dynamical Systems

### v1

@@ 4.2.2 Exploring evolution of connectivity over longer time periods
The third example, P3B1 (Figure 14), shows a growth and then stabilization of network complexity. At DIV 14, no active features are observed and most of the activity shown in subnetwork 0 is due to the low-level background process. From DIV 21-35, there is a developmental increase in the number of active features. We see high overlap in the active subnetworks: subnetwork $\theta_{1}$ persists from DIV 14 – DIV 35, and subnetworks $\theta_{2}$ and $\theta_{3}$ persist from DIV 28 to DIV 35.
@@ Conclusion
The model performs well on MEA data from murine primary hippocampal cultures and from human iPSC-derived cortical cultures. It captures a developmental increase in the number of subnetworks and the likelihood of these subnetworks being on during the MEA recordings. This model provides an important tool for neurobiologists studying brain development and brain disorders with MEA recordings. In extends the tools beyond pair-wise comparisons to detect functional connectivity and identifies patterns of activity from subnetworks of different sizes that can vary on the scale of seconds to minutes during the MEA recording. This will enable neurobiologists to investigate differences in network complexity, capacity for information sharing, and stability versus flexibility of state switching in subnetworks. As the method uses spike train data, it may also be useful for analysis of data from other multielectrode recordings (e.g., Neuropixel probes), which may be performed in vivo, although significantly more than 64 electrodes would likely increase the computational demands. This method could also be used with MEA recordings to compare the effect of drugs on modulating network states in vitro or the effect of genetic mutations on the number of states and frequency of state switching. Thus, this approach may inform future mechanistic and therapeutic investigations in human-derived or murine models of neurological disorders.

### v2

@@ 4.2.2 Exploring evolution of connectivity over longer time periods
The third example, P3B1 (Figure 5 of the Supplementary Material (Gong et al.,, 2025)), shows a growth and then stabilization of network complexity. At DIV 14, no active features are observed, and most of the activity shown in subnetwork 0 is due to the low-level background process. From DIV 21-35, there is a developmental increase in the number of active features. We see high overlap in the active subnetworks: subnetworks $\theta_{1}$ and $\theta_{2}$ persist from DIV21-28 and $\theta_{5}$ persists from DIV 28 – DIV 35. Additionally, three subnetworks dominated by a single electrode ($\theta_{6}$, $\theta_{7}$ and $\theta_{10}$) persist from DIV 14 to DIV 35.
In the bottom row of Table 1, we compare the average one-step-ahead predictive RMSE of FSLDS with rSLDS and LDS, on the three cultures explored in this section. For rSLDS, we select the number of latent states and latent features using held-out cross-validation on a forth culture P1A2. In this case, FSLDS does not perform better than rSLDS and LDS in terms of RMSE. However, this is perhaps not surprising: the held-out test set only contains timepoints in DIV35. Unlike the murine hippocampal data, in the iPSC cultures we do not see significant changes in the latent features within a time period, so there is no predictive benefit in modeling such transitions. However, this is offset by the fact that FSLDS does give us insights into the switching behavior between time periods, allowing insights about neuronal development that are not directly accessible via the existing LDS/rSDLS approaches.
@@ Conclusion
We proposed a new factorial switching dynamical system model to detect changes in brain networks. This model allows concurrent changes in subnetworks, without assuming unrealistic abrupt changes in the whole network. We pair it with a scalable inference algorithm enabling fast computation. The model performs well on MEA data from murine primary hippocampal cultures and from human iPSC-derived cortical cultures. It captures a developmental increase in the number of subnetworks and the likelihood of these subnetworks being on during the MEA recordings. Although LDS/SLDS occasionally achieves higher prediction accuracy, our model provides interpretable decomposition of network activity into constituent subnetworks and their temporal dynamics. Several natural extensions could be pursued in future research. For example, in this paper, we use a fixed number of features $K=10$ (in practice, we found ten features sufficient to model the real-world datasets explored here) and use a regularization parameter $\lambda$ to shrink the number of active features. An interesting future line of work would be to learn the number of features, either by imposing a prior on $K$ or by using a nonparametric Bayesian approach.
This model provides an important tool for neurobiologists studying brain development and brain disorders with MEA recordings. It extends the tools beyond pair-wise comparisons to detect functional connectivity and identifies patterns of activity from subnetworks of different sizes that can vary on the scale of seconds to minutes during the MEA recording. This will enable neurobiologists to investigate differences in network complexity, capacity for information sharing, and stability versus flexibility of state switching in subnetworks. As the method uses spike train data, it may also be useful for analysis of data from other multielectrode recordings (e.g., Neuropixel probes), which may be performed in vivo. This method could also be used with MEA recordings to compare the effect of drugs on modulating network states in vitro or the effect of genetic mutations on the number of states and frequency of state switching. Thus, this approach may inform future mechanistic and therapeutic investigations in human-derived or murine models of neurological disorders.